OpenStudy (anonymous):

the height of a triangle is 3 inches more than twice its base. If the area of the triangle is 16 square inches, find the base and the height of the triangle.

OpenStudy (nottim):

Use formula A=bh/2, where h=x+3, area=16 sq inches.

OpenStudy (anonymous):

i haves 2x+3-32=0 when i use the quadratic formula ... i am lost on how to divide the square root as in -3 - or + square root of 265 divided by 4

OpenStudy (nottim):

Er...I wouldn't know what you're talking about by the way you describe it.

OpenStudy (anonymous):

sorry its 2x^2 +3x -32

OpenStudy (nottim):

ok

OpenStudy (nottim):

Usually, you would not divide the square root to leave it an accurate answer.

OpenStudy (anonymous):

i have to find the base and height for a answer and my teacher wants it in a decimal

OpenStudy (nottim):

Ok. A calculator won't do?

OpenStudy (nottim):

Follow your PEDMAS/BEDMAS Rules then.

OpenStudy (anonymous):

thats the problem ... i dont know how exactly to put it on my calculator to solve it

OpenStudy (anonymous):

it comes up with a error

OpenStudy (nottim):

is there a minus sign INSIDE the sqr root?

OpenStudy (nottim):

are you using scientific calculator?

OpenStudy (anonymous):

my calculator is a TI-30XS

OpenStudy (nottim):

wazzat mean

OpenStudy (nottim):

yur lazer machine?

OpenStudy (nottim):

if it has a double line display, it should be one.

OpenStudy (anonymous):

its a scientific calculator

OpenStudy (nottim):

oh. remember to use brackets, if you're doing it in one step.

OpenStudy (nottim):

did you fall asleep typing?

OpenStudy (anonymous):

i dont understand how to divide this i have for my answer (-3-√(365))/(4) ... but its not done .. need to find the base and height the height of a triangle is 3 inches more than twice its base. If the area of the triangle is 16 square inches, find the base and the height of the triangle.

OpenStudy (anonymous):

dont know how to divide this on my calculator ...

OpenStudy (nottim):

uh...

OpenStudy (nottim):

Try entering it on wolfram alpha. the website.